# SoK: What don’t we know? Understanding Security Vulnerabilities in SNARKs

### v2

@@ Appendix B ZK Vulnerabilities Examples
[74] 2019
The trusted setup in Zcash’s Pinocchio-based implementation produced additional parameters that compromised the ’toxic waste’ secrecy. This vulnerability enabled malicious provers to create false proofs, leading to the unauthorized creation of Zcash tokens, effectively resulting in token counterfeiting.
@@
 [53] 2019
A vulnerability in circomlib’s implementation of the MiMC hash function, used in Tornado’s merkle tree of deposits, permitted the verification of incorrect hashes. This flaw could lead to unauthorized withdrawals of 0.1 ETH from Tornado.cash without legitimate deposits. The core issue was a computation in the hash function being performed without the necessary constraint in place, enabling this exploit.
Circuit Breaking Soundness
[90] 2021
The Aztec Plonk verifier had a flaw where setting two key elements to 0 caused it to mistakenly accept a proof, as it misinterpreted these zeros as the ‘point at infinity’ in elliptic curve cryptography. This vulnerability enables attackers to forge proofs by exploiting the verifier’s handling of these specific zero values.
Backend Breaking Soundness
 [23] 2022
In Solidity, the uint256 type’s capacity to exceed the snark scalar field order poses an overflow risk. To counter this, the Semaphore verifier smart contract fails if a public input, like a group ID, exceeds this order. If a group’s ID is set too high, it causes the verifier to revert, making the group non-functional. Hence, the Semaphore smart contract had to perform a check before passing data to the circuits.
Integration Breaking Completeness
@@
In Table 8, we present four examples of vulnerabilities relevant to our study on the security of SNARK systems. These examples demonstrate the range and complexity of possible security issues that could compromise the trustworthiness of systems based on SNARKs. To enhance our analysis and offer deeper understanding into these vulnerabilities, we further explore additional examples, including code snippets, focusing specifically on the circuit and integration layers.

### v4

@@ Appendix B ZK Vulnerabilities Examples
 [57] 2019 A vulnerability in circomlib’s implementation of the MiMC hash function, used in Tornado’s merkle tree of deposits, permitted the verification of incorrect hashes. This flaw could lead to unauthorized withdrawals of 0.1 ETH from Tornado.cash without legitimate deposits. The core issue was a computation in the hash function being performed without the necessary constraint in place, enabling this exploit. Circuit Breaking Soundness
[95] 2021 The Aztec Plonk verifier had a flaw where setting two key elements to 0 caused it to mistakenly accept a proof, as it misinterpreted these zeros as the ‘point at infinity’ in elliptic curve cryptography. This vulnerability enables attackers to forge proofs by exploiting the verifier’s handling of these specific zero values. Backend Breaking Soundness
 [26] 2022 In Solidity, the uint256 type’s capacity to exceed the snark scalar field order poses an overflow risk. To counter this, the Semaphore verifier smart contract fails if a public input, like a group ID, exceeds this order. If a group’s ID is set too high, it causes the verifier to revert, making the group non-functional. Hence, the Semaphore smart contract had to perform a check before passing data to the circuits. Integration Breaking Completeness
@@
In Table 8, we present four examples of vulnerabilities relevant to our study on the security of SNARK systems. These examples demonstrate the range and complexity of possible security issues that could compromise the trustworthiness of systems based on SNARKs. To enhance our analysis and offer a deeper understanding of these vulnerabilities, we further explore additional examples, including code snippets, focusing specifically on the circuit and integration layers.